OpenStudy (anonymous):

Twelve of the 40 cars pulled by the locomotive were tankers a.what fraction of the cars were tankers? b.what percent of the cars were not tankers?

OpenStudy (anonymous):

Very simple. Total vehicles = 40. Fraction of cars = cars / total = 12 / 40 = 30 percent or 3/10ths. Then, the percentage of vehicles that were not tankers is 100% - 30% or 1 - 0.3 = 70% or 0.7.